A sample of the black mineral hematite, an oxide of iron found in many iron ores. contains 34.97 g of iron and 15.03 g of oxygen . What is the empirical formula of hematite

1
Expert's answer
2021-10-30T01:13:10-0400

Fe = 34.97g

O = 15.03g


Divide by molar mass

Fe = 34.97/56 = 0.624

O = 15.03/16 = 0.939


Divide by smaller number

Fe = 0.624/0.624 = 1

O = 0.939/0.624 = 1.5


Multiply by 2 to turn values to natural numbers

Fe = 2

O = 3


"\therefore" The empirical formula of the compound is "Fe_2O_3"